(P = 0) 0 0 Fia. Q2.

2. A helical compression spring is to be designed as shown in Fig Q2. The spring is required to exert a force of Po-102 N when compressed to a length of lo-44 mm. When its length is I-76 mm, it must exert a force of P-22 N. The spring will be cycled rapidly, with severe service required. Use ASTM A401 steel wire. The factor of safety, η, need to be larger than 1.5. Starting with a spring index of 9, and checking for stability is required. a) Compute the spring rate k and the free length Ir of the spring. (6 pts) b) Determine the wire diameter dw, of the spring to satisfy the conditions given, a a peened spring. Use the Goodman failure locus for the Zimmerli data. (11 pts) termine the solid length l, the force P, and stress corresponding to the solid length, the number of active coils, OD of the coil and the factor of safety guarding against statioc ssuming failure. (18 pts)
(P = 0) 0 0 Fia. Q2.
Answer #1

From the required force condition,

when Pr = 102 N, Li = 76 mm and Po = 22 N, Lo = 44 mm

As we know force by the spring is directly proportional to the deflection of the spring form free length i. e. Lf.

Hence Pr = k (Lf - Li)

Po = k (Lf - Lo)

Putting values

102 = k (Lf - 76) .... (1)

22 = k (Lf - 44) .... (2)

Solving equation (1) and (2)

we get free length Lf = 35.2 mm and spring constant k = 2.5 N/mm .... Ans

From property data sheet of ASTM A401 steel,

Ultimate tensile strength Sut = 1050 N/mm2

Modulus of rigidity G = 81370 N/mm2

Required FOS = 1.5 and from Goodman failure condition;

FOS

where Ssy is torsional yield strength

Ssy = 0.577 Syt where Syt is tensile yield strength

and Syt = 0.75 Sut

т=0.3 * Sut

τ = 315N/mm2

Spring index C = 9

Wahl factor is given as

4C-1 4C-4

Hence K = 1.093

Shear stress is give by

8 PC

8 102 9 315 = 1.093 *

\thereforeWire diameter dw = 2.848 mm ..... Ans

Mean coil diameter D = C * dw = 9 * 2.848 = 25.63 mm

Number of active coils N is given by below formula

\delta = \frac{8 * P * D^{3} * N}{G * d_{w}^{4}}

where \delta is spring deflection for force P

\delta = 22/2.5 = 8.8 mm

N is number of active coils

8 * N 22 * 25633 813702.844

\thereforeNumber of active coils N = 15.72 ~ 16 .... Ans

Assuming number of inactive coils = 2 (for peened spring)

Total coils Nt = N + 2 = 18

Solid length of the Spring = Nt * dw = 18 * 2.84 = 51.12 mm .... Ans

Force required for compressing spring to solid length Ps = K * solid length

Ps = 51.12 * 2.5 = 127.8 N ..... Ans

Stress corresponding to solid length

\tau = K *\frac{8*P_{s}*C}{\pi * d_{w}^{2}}

\tau = 1.093 *\frac{8* 127.8*9}{\pi * 2.84^{2}}

\tau = 396.91 N/mm2 ..... Ans

Outer diameter of spring Do = D + dw/2

Do = 25.63 + (2.84)/2 = 27.05 mm .... Ans